### v1

# Point configurations in sets of sufficient topological structure and a topological Erdős similarity conjecture

Abstract.
We explore the occurrence of point configurations within non-meager (second category) Baire sets. A celebrated result of Steinhaus asserts that $A+B$ and $A-B$ contain an interval whenever $A$ and $B$ are sets of positive Lebesgue measure in $\mathbb{R}^{n}$ for $n\geq 1$. A topological analogue attributed to Piccard asserts that both $AB$ and $AB^{-1}$ contain an interval when $A,B$ are non-meager (second category) Baire sets in a topological group. We explore generalizations of Piccard’s result to more complex point configurations and more abstract spaces. In the Euclidean setting, we show that if $A\subset\mathbb{R}^{d}$ is a non-meager Baire set and $F=\{x_{n}\}_{n\in\mathbb{N}}$ is a bounded sequence, then there is an interval of scalings $t$ for which $tF+z\subset A$ for some $z\in\mathbb{R}^{d}$. That is, the set
$\Delta_{F}(A)=\{t\in\mathbb{R}:\exists z\text{ such that }tF+z\subset A\}$
has nonempty interior. More generally, if $V$ is a topological vector space and $F=\{x_{n}\}_{n\in\mathbb{N}}\subset V$ is a bounded sequence, we show that if $A\subset V$ is non-meager and Baire, then $\Delta_{F}(A)$ has nonempty interior. The notion of boundedness in this context is described below. Note that the sequence $F$ can be countably infinite, which distinguishes this result from its measure-theoretic analogue. In the context of the topological version of Erdős’ similarity conjecture, we show that bounded countable sets are universal in non-meager Baire sets.
1. Introduction
A large body of research concerns the occurrence (or absence) of patterns in sets of sufficient Hausdorff dimension, see for instance, [5, 6, 7, 15, 19], or in sets of sufficient Fourier dimension [3, 1, 24]. Another notion of size that has proved useful in the study of point configurations is that of Newhouse thickness, see for instance, [16, 17, 27, 23]. In this article, we focus on the occurrence of infinite point configurations within sets of sufficient topological structure.
Our story begins with a well-known theorem of Steinhaus from the 1920s and its topological analogue, known as Piccard’s theorem. Let $A,B\subset\mathbb{R}^{n}$ denote sets of positive measure. Steinhaus’ theorem states that the set $A-B$ contains a nontrivial open set (see [26] for $n=1$ and [12] for $n\geq 1$). Said differently, there is an open set of shifts $t$ so that
(1) $A\cap(B+t)\neq\varnothing.$
This assertion can be proved by observing that the convolution of the characteristic functions $\chi_{A}$ and $\chi_{B}$ is a continuous function, so that the function
[…]
is continuous and, by Fubini, not identically zero. Alternatively, (1) follows as a simple consequence of the Lebesgue density theorem.
For more intricate configurations, these same proof techniques can be recycled to show that for each $k\in\mathbb{N}$, there exists an open ball $S_{k}$ about the origin so that if $\{v_{i}\}_{i=1}^{k}\subset S_{k}$, then
(2) $\bigcap_{i=1}^{k}(A+v_{i})\neq\varnothing.$
It is an immediate consequence that, if $F=\{v_{i}\}_{i=1}^{k}$ denotes an arbitrary finite point configuration, then $A$ contains all sufficiently small affine copies of $F$. That is, if
$\Delta_{F}(A)=\{t\in\mathbb{R}:\exists z\in\mathbb{R}^{n}\ tF+z\subset A\},$
then $\Delta_{F}(A)$ contains an interval with left endpoint 00.
We note that there are many further generalizations and modifications of Steinhaus’ theorem, including a result of Erdős and Oxtoby showing that if $F$ on $\mathbb{R}\times\mathbb{R}$ is a continuously differentiable real-valued function with non-vanishing partial derivatives on its domain, then $F(A,B)$ has nonempty interior [2]. This theorem also holds in higher dimensions, see [8, Chapter II] and the references therein and [25, Proposition 2.9 (i)].
A topological analogue of Steinhaus’ theorem known as Piccard’s theorem asserts that the sum of two non-meager Baire sets (see Definition 1.1) in a topological vector space contains a nonempty open set [21, 13]. Again, addition can be replaced by a two variable function satisfying minimal differentiability assumptions, see [25, Proposition 2.9 (iii)] and [8, Chapter II]. Also see [22] for an analogous statement for topological groups, which states that $A*B$ (and $A*B^{-1}$) contains a nonempty open set provided that $A$ and $B$ are non-meager Baire sets.
While the aforementioned results consider operations on pairs of points, the present article focuses on general point configurations. This work fits into a large and expanding literature on the presence of geometric patterns in fractal sets. The basic problem is to understand how large a subset of Euclidean space must be to ensure that it contains the vertices of a congruent and possibly scaled copy of a given polyhedron or another geometric shape.
This article fits into the broader context of investigating a topological variant of the Erdős similarity conjecture:
For each uncountable set $F$, there exists a non-meager Baire set $A$ that contains no affine image of $F$.
We briefly summarize what is known about this conjecture. Let $E\subset\mathbb{R}^{d}$, and let $X$ denote a collection of subsets in $\mathbb{R}^{d}$. We say that $E$ is universal in $X$ if every $K\in X$ contains an affine copy of $E$.
Universality can be considered for different classes of sets. In the setting of the classic Erdős similarity conjecture, there is a distinction between considering universality over full measure sets versus universality over measurable sets. This distinction persists for the topological Erdős similarity conjecture with co-meager sets and non-meager sets in place of full measure and positive measure respectively. An example of a set that is non-meager but not co-meager is $[0,1]$; more generally, any set with both nonempty interior and exterior is a non-meager set which is not co-meager. For the full story on the Erdős similarity conjecture, see [10].
It is a straightforward consequence of the Baire category theorem that countable sets are universal in the collection of dense $G_{\delta}$ sets (see Example A.9 in Section A). Further, since every co-meager set contains a dense $G_{\delta}$ set, it follows that the countable sets are universal in the collection of co-meager sets.
Our main Euclidean result (Theorem 2.1) extends this observation to non-meager Baire sets; we show that bounded countable sets are universal in non-meager Baire sets. Further, we prove a more quantitative result about an interval worth of scalings, and our results extend to topological vector spaces (Theorem 3.3).
For uncountable sets, Gallagher, Lai, and Weber [4] proved that certain Cantor sets are not universal in non-meager Baire sets, so that Theorem 2.1 need not generically hold with countable replaced by uncountable (see Example A.8 in Section A. Also see, [9] where Jung and Lai demonstrate that a set is universal in the collection of dense $G_{\delta}$ sets if and only if it is a strong measure zero set.
1.1. Background and definitions
Before stating our main results, we give some definitions and background results.
Definitions 1.1.
Let $X$ denote a topological space. A subset of $X$ is meager (first category) if it is a countable union of nowhere dense sets. A set is non-meager (second category) if it is not meager. Finally, a set $A\subset X$ is said to be Baire if there is an open set $U$ and a meager set $P$ so that $A$ equals the symmetric difference $A=U\Delta P$.
Remark 1.2.
Equivalently, one may define Baire sets as sets $A$ for which there exists an open set $U$ such that $U\Delta A$ is meager. If $A$ satisfies this condition, then $A$ satisfies Definition 1.1 with $P=U\Delta A$. Conversely, if $U$ satisfies Definition 1.1, then $A\Delta U=P$ is meager. See Theorem A.1 for other useful equivalent definitions.
Morally, one may think of meager sets as the topological analogues of measure zero sets (in the sense that they are considered negligible and are stable under countable unions), and of Baire sets as the topological analogues of measurable sets (in the sense that they are “almost” equal to open sets up to negligible error). A word of caution: even meager sets can have full measure. A succinct account of basic facts and examples is given in Appendix A; for a more thorough introduction, see [11, Chapter I.8].
Theorem 1.3 (Piccard - Euclidean setting).
If $A,B\subset\mathbb{R}^{d}$ are non-meager Baire sets, then both $A+B$ and $A-B$ have nonempty interior.
A short proof of Theorem 1.3 is given in Section 2 for emphasis. However, we also give a proof of the following more general result in Section 3.
Theorem 1.4 (Piccard - Abstract version).
Let $X$ denote a topological vector space (or a topological group). If $A,B\subset X$ are non-meager Baire sets, then both $A+B$ and $A-B$ (both $A*B$ and $A*B^{-1}$) contain nonempty open sets.
The definition of $\Delta_{F}(A)$ given above immediately extends to vector spaces:
Definition 1.5.
Let $V$ be a real vector space. Given sets $A,F\subset V$, define
$\Delta_{F}(A)=\{t\in\mathbb{R}:\exists z\in V\ tF+z\subset A\}.$
In the notation above, a consequence of Piccard’s Theorem is that $\Delta_{F}(A)$ has nonempty interior whenever $A$ is a non-meager Baire set and $F$ is a two-point set. In the following sections, we explore generalizations of this result to more complex point configurations and more abstract spaces. To make sense of the set $\Delta_{F}(A)$, we need both translations and scalings, and for this reason we work in topological vector spaces.
1.2. Acknowledgment
K.T. is supported in part by the Simons Foundation Grant GR137264. Both authors thank Angel Cruz, Yeonwook Jung, and Yuveshen Mooroogen for interesting discussions related to this article.
2. Results in Euclidean space
Our first main result is the following generalization of Piccard’s Theorem.
Theorem 2.1.
Let $A\subset\mathbb{R}^{d}$ be a non-meager Baire set, and let $F=\{x_{n}\}_{n\in\mathbb{N}}$ be a bounded sequence. Then, $\Delta_{F}(A)$ has nonempty interior.
The necessity of the hypotheses are exemplified in Section A. The proof of Theorem 2.1 is a topological variant of the classic proof of Steinhaus’s Theorem. The main idea of Steinhaus’s Theorem via the Lebesgue density theorem is as follows. If $A\subset\mathbb{R}^{d}$ is a measurable set, then a ball $B$ can be chosen so that $A$ takes up $99\%$ of the space of $B$. This means $A$ is guaranteed to intersect small translations of itself, which is the crux of the proof of the theorem. This approach works for finite configurations; as there is finite loss of measure with each intersection. If $A$ is instead a non-meager Baire set, we may still approximate $A$ by a ball in the following sense. The following is immediate from the definition of a Baire set.
Lemma 2.2 (Topological variant of Lebesgue Density Theorem).
Let $A\subset\mathbb{R}^{d}$ be a non-meager Baire set. There exists an open ball $B$ such that $B\setminus A$ is meager.
Proof.
Since $A$ is Baire, there is an open set $U$ such that $A\Delta U$ is meager (since $A=U\Delta P$ implies $A\Delta U\subset P$). Since $A$ is non-meager, $U$ must be nonempty. Let $B$ be any open ball contained in $U$. Since
$(B\setminus A)\subset(U\setminus A)\subset A\Delta U,$
the set $B\setminus A$ is meager. ∎
Recall that in the Baire Category setting, we think of meager sets as the analogues of measure-zero sets in measure theory; that is, this is the relevant notion of a set being negligible, and this notion is stable under countable unions. However, the Lebesgue Density Theorem only guarantees that a ball can approximate a measurable set with (say) a $1\%$ error. On the other hand, Lemma 2.2 says that a ball may approximate $A$ with meager error. This stronger approximation lemma is the reason we can find countable configurations, rather than just finite ones.
Before we prove our main theorem, we give a short proof of Piccard’s Theorem using Lemma 2.2 in the case $d=1$ (the proof when $d\geq 2$ is a simple exercise). Theorem 1.3 of course follows from Theorem 1.4, but we give the proof here anyways explicitly in the Euclidean setting.
Proof of Theorem 1.3 (one dimensional case).
For $t\in\mathbb{R}$, define $B_{t}:=t-B$. We have $t\in A+B$ if and only if $A\cap B_{t}\neq\varnothing$. Apply Lemma 2.2 to get non-degenerate intervals $I,J$ such that $I\setminus A$ and $J\setminus B$ are meager, and let $J_{t}=t-J$. For any $t$, the set $J_{t}\setminus B_{t}$ must be meager also. Therefore, the set
$(I\cap J_{t})\setminus(A\cap B_{t})\subset(I\setminus A)\cup(J_{t}\setminus B_{t})$
is meager. If $A\cap B_{t}$ were meager, it would follow that $(I\cap J_{t})$ is meager. This cannot happen if $t$ is chosen so that $I\cap J_{t}$ is a non-degenerate interval. Therefore, for an interval worth of $t$, we have $A\cap B_{t}\neq\varnothing$. Equivalently, $A+B$ contains an interval. ∎
This proof idea generalizes to arbitrary countable configurations without too much difficulty. The key geometric fact we need to use is that the intersection of infinitely many balls still has nonempty interior, provided that each ball is a small translate of a single original ball.
Proof of Theorem 2.1.
Define $A_{n,t}=A-tx_{n}$. We have $\{tx_{n}+z:n\in\mathbb{N}\}\subset A$ if and only if $z\in\bigcap_{n=1}^{\infty}A_{n,t}$. We must show this intersection is nonempty for an interval worth of $t$. Apply Lemma 2.2 to obtain an open ball $B$ with $B\setminus A$ meager, and let $B_{n,t}=B-tx_{n}$. It follows that $B_{n,t}\setminus A_{n,t}$ is meager for every $n,t$. Finally, let
$\widetilde{B}_{t}=\bigcap_{n=1}^{\infty}B_{n,t}.$
The set $\widetilde{B}_{t}\setminus A_{n,t}$ is meager, since $\widetilde{B}_{t}\subset B_{n,t}$. Therefore,
$\widetilde{B}_{t}\setminus\left(\bigcap_{n=1}^{\infty}A_{n,t}\right)=\bigcup_{n=1}^{\infty}(\widetilde{B}_{t}\setminus A_{n,t})$
is meager. If $\bigcap_{n=1}^{\infty}A_{n,t}$ were meager, then $\widetilde{B}_{t}$ would be meager. However, if $t$ is sufficiently small (relative to the radius of $B$ and $\sup_{n}|x_{n}|$), then each set $B_{n,t}$ is a small translate of the original ball $B$, and hence the intersection $\widetilde{B}_{t}$ has nonempty interior. Therefore, the theorem holds when $J$ is a sufficiently small neighborhood of 00. ∎
3. Results in topological vector spaces
The purpose of this section is to generalize our result in Euclidean space (Theorem 2.1) to the more general setting of topological vector spaces. We start by making some definitions.
Definition 3.1.
A topological vector space is a vector space $V$ with a topology such that the operations of addition and scalar multiplication are continuous as functions $V\times V\to V$ and ×$\mathbb{R}\times V\to V$, respectively. We will denote the zero vector by a bold 0, to contrast with the zero scalar 00.
As a warm-up, we give the proof of Theorem 1.4. The following proof also works in the case that $X$ is a topological group; one only need change the notation and replace all occurrences of $A-B$ with $A*B^{-1}$.
Proof of Theorem 1.4.
We establish the claim for $A-B$. Let $U_{1}$, $U_{2}$ denote open sets, and let $P_{1}$, $P_{2}$ denote meager sets so that $A=U_{1}\Delta P_{1}$ and $B=U_{2}\Delta P_{2}$. Let us assume that nonempty open sets are not meager, for otherwise $X$ would also be meager by the Banach category theorem (see Theorem 3.4) as a union of translates of such a set.
We observe that
(3) $A-B\supset U_{1}-U_{2},$
which will prove the claim since $U_{1}-U_{2}$ is a nonempty open set. To show (3), let $t\in U_{1}-U_{2}$. Observe that $t\in A-B$ if and only if $\left(t+B\right)\cap A\neq\emptyset.$ Since
(4) $A\cap\left(t+B\right)\supset(U_{1}\cap(t+U_{2}))\backslash(P_{1}\cup(P_{2}+t)),$
and since $(U_{1}\cap(t+U_{2}))$ is a nonempty open set, while $(P_{1}\cup(P_{2}+t))$ is meager, it follows that the right-hand-side of (4) is nonempty.
Moving on, recall that in Theorem 2.1, our hypotheses included that the configuration under consideration was bounded. Since an arbitrary topological vector space need not be metric, we need a definition of boundedness in this context.
Definition 3.2.
Let $V$ be a topological vector space, and let $F\subset V$. We say $F$ is bounded if, for any open set $U\subset V$, there exist $t>0,z\in V$ such that $tF+z\subset U$.
We are now ready to state the main theorem of this section.
Theorem 3.3.
Let $V$ be a topological vector space, and let $F=\{x_{n}\}_{n}\subset V$ be a bounded sequence. If $A\subset V$ is a non-meager Baire set, then $\Delta_{F}(A)$ has nonempty interior.
The proof of Theorem 2.1 depended on several topological properties of the space $\mathbb{R}^{d}$, and we need analogues of those properties to allow the proof to be carried out in this setting. First and most obviously, our arguments in the Euclidean setting are based on the fact that $\mathbb{R}^{d}$ is a Baire space, which is a consequence of the Baire Category Theorem, so that nonempty open sets are non-meager. In the more general context of topological vector spaces, the Banach Category Theorem plays this role.
Theorem 3.4 (Banach Category Theorem).
In any topological space, the union of any family of meager open sets is meager.
See [20, Chapter 16] for an exposition of this result. As an immediate consequence, we have:
Lemma 3.5.
If $V$ is a topological vector space containing a non-meager set, then any nonempty open subset of $V$ must be non-meager.
Proof.
Let $V$ be such. If there were a meager nonempty open subset $U$ of $V$, then $V$ would itself be meager by the Banach category theorem as the union of translates of $U$. But if $V$ were meager, then it could not contain a non-meager subset. ∎
Next, we need a topological vector space analogue of the open balls of Euclidean space. In Definition 3.2, if $F$ is a bounded set and we scale $F$ by $t_{0}$ to be in a given neighborhood of 0, then there is no guarantee that we could then scale by a smaller $t<t_{0}$ and still guarantee this containment. Fortunately, in any topological vector space, there is a topological base consisting of sets which are closed under scaling down. This is captured in the following lemma.
Definition 3.6.
Let $V$ be a topological vector space. A set $B$ is called balanced if for any $t\leq 1$, we have $tB\subset B$.
Lemma 3.7 ([18], Theorem 4.5.1).
Let $V$ be a topological vector space. There is a neighborhood base at 0 consisting of balanced open sets.
In the proof of Theorem 2.1, the way boundedness is used is that a bounded set multiplied by a small scalar produces a set which is contained in a small ball around the origin. We then observe that an arbitrary intersection of open sets is guaranteed to be open, provided each of the sets is a small perturbation of a fixed interval. The following lemma will play the analogous role in this section and is a key ingredient in the proof of our main result.
Lemma 3.8.
Let $V$ be a topological vector space, let $U$ be an open set, and let $F$ be a bounded set. For each $x\in F$, define $U_{x,t}=U-tx$. Finally, define
$\widetilde{U}_{t}=\bigcap_{x\in F}U_{x,t}.$
There exists a non-degenerate interval $J\subset\mathbb{R}$ such that for any $t\in J$, the set $\widetilde{U}_{t}$ has nonempty interior.
Proof.
First, we consider the case $\textbf{0}\in U$. Since vector space addition $+:V\times V\to V$ is continuous, the pre-image set
$\{(x,y)\in V\times V:x+y\in U\}$
is open. Since Cartesian products of open sets are a basis for the product topology, there exist open sets $U_{1},U_{2}\subset V$ such that $(\textbf{0},\textbf{0})\in U_{1}\times U_{2}$ and $U_{1}+U_{2}\subset U$. The set $U^{\prime}=U_{1}\cap U_{2}\cap U$ is therefore an open neighborhood of 0, contained in $U$, with the property that $U^{\prime}+U^{\prime}\subset U$. By Lemma 3.7, we may take a subset $B\subset U^{\prime}$ with these same properties, which is also balanced. Since $F$ is bounded, there exists $t_{0}$ such that $t_{0}F\subset B$. Since $B$ is balanced, we therefore have $tF\subset B$ for any $t\in J:=(0,t_{0})$. We claim that for any such $t$, the set $\widetilde{U}_{t}$ contains $B$ as a subset. Indeed, for any $z\in B,t\in J$, and $x\in F$, we have
$z+tx\in B+B\subset U,$
hence
$z=(z+tx)-tx\in U-tx=U_{x,t}.$
Now, consider the general case. For any open set $U$, we consider a translation $U^{\prime}$ of $U$ that contains the origin. From the first case, it follows that $\widetilde{U^{\prime}}_{t}$ has a nonempty interior for an interval worth $t$. Since $\widetilde{U}_{t}$ is a translate of $\widetilde{U^{\prime}}_{t}$, the result follows. ∎
We are now ready to prove the main theorem of this section. For variety of exposition, rather than just present a modification of the proof of Theorem 1.4 given at the beginning of this section, we present an alternative proof.
Proof of Theorem 3.3.
Define $A_{n,t}=A-tx_{n}$. We have $\{tx_{n}+z:n\in\mathbb{N}\}\subset A$ if and only if $z\in\bigcap_{n=1}^{\infty}A_{n,t}$. We must show that this intersection is nonempty for an interval worth of $t$. Let $U$ be an open set and let $P$ be a meager set such that $A=U\Delta P$; let $U_{n,t}=U-tx_{n}$ and $P_{n,t}=P-tx_{n}$. Since any translation of a nowhere dense set is still nowhere dense, the set $P_{n,t}$ is meager for every $n,t$. Define
$\widetilde{U}_{t}=\bigcap_{n=1}^{\infty}U_{n,t}\,\,\,\text{ and }\,\,\,\widetilde{P}_{t}=\bigcup_{n=1}^{\infty}P_{n,t}.$
We now observe that
$\bigcap_{n=1}^{\infty}A_{n,t}\supset\widetilde{U}_{t}\backslash\widetilde{P}_{t}.$
The proof is completed by arguing that $\widetilde{U}_{t}$ is non-meager; indeed, since $\widetilde{P}_{t}$ is meager, then it would follow that $\bigcap_{n=1}^{\infty}A_{n,t}$ is nonempty. By Lemma 3.8, there is a non-degenerate interval $J\subset\mathbb{R}$ such that for any $t\in J$, the set $\widetilde{U}_{t}$ has nonempty interior. By Lemma 3.5, since $V$ contains a non-meager set $A$, a set with nonempty interior must be non-meager.
Appendix A Examples and properties of Baire sets
The purpose of this appendix is to demystify the definition of a Baire set and describe some examples which show that the assumptions of our main theorems cannot be dropped. First, we have the following alternate characterizations of Baire sets.
Theorem A.1 ([11], Proposition 8.23).
Let $X$ be any topological space, and let $A\subset X$. The following are equivalent:
  1. (i)
    $A$ is a Baire set.
  2. (ii)
    $A=G\cup P$ for some $G_{\delta}$ set $G$ and some meager set $P$.
  3. (iii)
    $A=F\setminus P$ for some $F_{\sigma}$ set $F$ and some meager set $P$.
As an immediate consequence, we have the following classes of examples of Baire sets.
Corollary A.2.
Let $X$ be any topological space.
  1. (a)
    Every meager subset of $X$ is Baire.
  2. (b)
    Every $G_{\delta}$ or $F_{\sigma}$ subset of $X$ is Baire.
It turns out that these examples generate all possible Baire sets. More precisely, we have.
Theorem A.3 ([11], Proposition 8.22).
In any topological space, the collection of Baire sets is the smallest $\sigma$-algebra containing all open sets and all meager sets.
Since the class of Borel sets is defined as the smallest $\sigma$-algebra containing all open sets, we have:
Corollary A.4.
All Borel sets are Baire.
The most well-known example of a non-measurable set is also an example of a non-Baire set. It also shows the necessity of the assumption in our results that our sets are Baire.
Example A.5 (A non-meager set for which the conclusion of Theorem 2.1 fails).
Let $\sim$ denote the equivalence relation on $\mathbb{R}$ given by $x\sim y$ if and only if $x-y\in\mathbb{Q}$. Let $A$ be a set consisting of exactly one representative of each equivalence class. We claim $A$ is non-meager; indeed, if $A$ were meager, then every translate of $A$ would also be meager, and thus
$\mathbb{R}=\bigcup_{q\in\mathbb{Q}}(A+q)$
would be meager (this is pointed out in [20], where it is used to prove directly that $A$ is not Baire). Let $F=\{x,y\}\subset\mathbb{R}$, where $x-y\in\mathbb{Q}$. By definition, for any $t\in\mathbb{Q}$ and $z\in\mathbb{R}$ we have $(tx+z)\sim(ty+z)$. By construction, $A$ cannot contain two distinct equivalent numbers. It follows that $\Delta_{F}(A)$ does not contain any rational number, and hence cannot have nonempty interior (in light of Theorem 2.1, this gives a less-direct proof that $A$ is not Baire).
There is a broad literature on fractal sets which avoid given configurations, and generally, such a set must be meager. This makes it very easy to import examples which show that Theorem 2.1 (and hence the more general Theorem 3.3) fail if a set is assumed to be Baire but not necessarily non-meager. For example, for any dimension $d$, there are well-known constructions of compact sets $A$ (of full dimension) such that $A-A$ has empty interior. In the language of this paper, this is the same as saying $\Delta_{F}(A)$ has empty interior for any set $F$ with at least two points. Beyond this, we can give examples of $3$-point sets $F$ and Baire sets $A$ of positive Hausdorff dimension where the set $\Delta_{F}(A)$ is actually empty (rather than simply having empty interior). To do this, we use a construction of Mathe.
Theorem A.6 ([14]).
Let $P(X_{1},\dots,X_{m})$ be a polynomial with rational coefficients in $dm$ variables (where each $X_{i}$ represents a vector variable in $\mathbb{R}^{d}$). There exists a compact set $E\subset\mathbb{R}^{d}$ of positive Hausdorff dimension which does not contain any configuration of $m$ distinct points $x_{1},\dots,x_{m}\in\mathbb{R}^{d}$ which lie in the zero set of $P$.
Example A.7 (An meager set with positive Hausdorff dimension for which the conclusion of Theorem 2.1 fails).
Let $x,y,z\in\mathbb{R}^{d}$ be three distinct points such that the number
$\alpha:=\frac{|y-x|}{|z-x|}$
is algebraic. It follows that $\alpha^{2}$ is also algebraic, so we may take
$Q(X):=a_{n}X^{n}+a_{n-1}X^{n-1}+\cdots+a_{1}X+a_{0}$
to be the minimal polynomial of $\alpha^{2}$ (here, $X$ is a one-dimensional variable rather than a vector variable as in Theorem A.6). Plugging in our definition of $\alpha$, we have
$a_{n}\left(\frac{|y-x|^{2}}{|z-x|^{2}}\right)^{n}+a_{n-1}\left(\frac{|y-x|^{2}}{|z-x|^{2}}\right)^{n-1}+\cdots+a_{1}\left(\frac{|y-x|^{2}}{|z-x|^{2}}\right)+a_{0}=0.$
Finally, multiplying through by $|z-x|^{2n}$ to clear denominators gives us
$a_{n}|y-x|^{2m}+a_{n-1}|y-x|^{2n-2}|z-x|^{2}+\cdots+a_{1}|y-x|^{2}|z-x|^{2n-2}+a_{0}|z-x|^{2n}=0.$
Thus, the points of $F$ satisfy a polynomial equation with rational coefficients. Moreover, by homogeneity and translation invariance, one can easily verify that the equation remains satisfied if $x,y,z$ are replaced by $tx+w,ty+w,tz+w$, respectively, for any $t>0$ and $w\in\mathbb{R}^{d}$. By Theorem A.6, it follows that there exists a compact set $A\subset\mathbb{R}^{d}$ such that $\Delta_{F}(A)=\varnothing$. Such a set $A$ must have empty interior, so it is nowhere dense and hence meager. However, because it is compact, it is still a Baire set by Corollary A.2. This shows the non-meager assumption in our theorems is necessary.
Our next example shows that the assumption that $F$ is countable in Theorem 2.1 is necessary.
Example A.8 (A non-meager Baire set for which Theorem 2.1 fails if countable is replaced by uncountable).
As metioned in the introduction, Gallagher, Lai, and Weber [4, Theorem 1.5] showed that Cantor sets on $\mathbb{R}$ with positive Newhouse thickness are not universal in the set of dense $G_{\delta}$ sets. In particular, they showed that there exists a dense $G_{\delta}$ set $G$ of full measure such that, for all Cantor sets $C$ with positive Newhouse thickness, $G$ does not contain an affine copy of $C$. The classic middle-thirds Cantor set, for example, has positive Newhouse thickness, and hence is not universal in the set of dense $G_{\delta}$ sets. Note that their $G$ is Baire, since it is a $G_{\delta}$ set, and $G$ is nonmeager since every dense $G_{\delta}$ set is comeager and hence non-meager.
Now, we consider an example that shows that our main Euclidean result, Theorem 2.1, is immediate for a special class of non-meager sets, mainly co-meager sets.
Example A.9 (countable sets are universal in co-meager sets).
Suppose that $G$ is meager and $F$ is countable. Observe that the Minkowski sum $G+F$ is meager as this set can be expressed as the countable union of shifts of the meager set $G$. Since $\mathbb{R}$ is not meager (by the Baire category theorem), then $\mathbb{R}\neq G+F$. Similarly, $\mathbb{R}\neq G-F$. This is equivalent to $(x+F)\cap G=\emptyset$ for some $x\in\mathbb{R}$. It follows that, if $F$ is countable and $C$ is co-meager, then $F+x\subset C$ for some $x\in\mathbb{R}$.
Our final (classic) example shows that although measure zero sets and meager sets can both be considered “negligible”, these two notions of negligibility are not mutually consistent.
Example A.10 (A meager set with full measure; a second category set with zero measure).
Enumerate the rationals by $q_{1},q_{2},\cdots$, and let $O(i,j)$ denote the open interval with center $q_{j}$ and length $\frac{1}{2^{i+j}}$. Set $U_{i}=\bigcup_{j=1}^{\infty}O(i,j)$, and set $D=\bigcap_{i=1}^{\infty}U_{i}$. One may verify that $D$ is a set of measure zero, while $\𝐷M=\mathbb{R}\backslash D$ is meager (as the complement of an intersection of open dense sets; each $U_{i}$ contains $\mathbb{Q}$). In conclusion, $\mathbb{R}=M\bigcup D$ can be expressed as the disjoint union of a meager set with infinite measure and a co-meager set (second category) with zero measure.